Given the intensity of an earthquake wave traveling through the Earth and detected 200 km from the source is 2 MW/m^2. What is the intensity of that wave if detected 300 km from the source?
1
Expert's answer
2021-07-19T09:52:04-0400

"I\sim \frac1{r^2},"

"I_2=I_1\frac{r_1^2}{r_2^2}=2\cdot\frac{200^2}{300^2}=0.89~\frac{MW}{m^2}."